My Wonderful Article in IEEE Format

I. M. Author  M. Y. Coauthor
My Department  Coauthor Department
My Institute  Coauthor Institute
City, ST zipcode  City, ST zipcode

Abstract

This is the abstract of my paper. It must fit within the size allowed, which is about 3 inches, including section title, which is 11 point bold font. If you don’t want the text in italics, simply remove the ’em’ command and the curly braces which bound the abstract text. If you have em commands within an italicized abstract, the text will come out as normal (nonitalicized) text.

1 Introduction

Here is my introduction text. There are 2 blank lines before the section heading and one afterwards. Heading text is 12 point bold font. Paragraphs are not indented. I may want a numbered subsection, which is done as follows.

1.1 Previous Work

In subsections there is 1 blank line before the section heading and one afterwards. Heading text is 11 point bold font. Paragraphs are separated by blank lines without tabs.

Throughout I may cite references of the form [1] or [2], and LaTeX will keep track of numbering. The numbers are based on the order you place them in the bibliography, not the order they appear in the text. They should (I believe) be in alphabetical order. La-Tex will put square brackets about the number within the text of your paper. For those of you new to the bibliography package, you may have to run the latex process twice to allow all references to be resolved. You will get a warning about a missing .aux file. Just rerun latex and it will be ok.

2 Summary and Conclusions

This template will get you through the minimum article, i.e. no figures or equations. To include those, please refer to your LaTeX manual and the IEEE publications guidelines. Good Luck!

Acknowledgements

This is how to do an unnumbered subsection, which comes out in 11 point bold font. Here I thank my colleagues, especially Mike Gennert, who know more about Tex and LaTeX than I.

References

[1] I. M. Author, “Some Related Article I Wrote,” Some Fine Journal, Vol. 17, pp. 1-100, 1987.

[2] A. N. Expert, A Book He Wrote, His Publisher, 1989.